# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ 4.1 Weak core
We use this example in our gadgets to prove that deciding the emptiness of the weak core is hard.
@@
It is NP-hard to decide whether a matching in the weak core exists, even if each partition class has size at most 7.
Before proving the theorem, we introduce the notion of special edges, which are actually not edges but gadgets that simulate edges with some special property. We say that there is a special edge between two vertices $u$ and $v$ if there is a gadget described below between $u$ and $v$. Let $H$ be a copy of the instance in Example 20. We have a player $S$ with four vertices $s_{1},s_{2},s_{3}$ and $s_{4}$. There is an edge between $s_{2}$ and $a_{1}\in H$. We also have a player $T$ with two vertices, $t_{1}$ and $t_{2}$ and edges $s_{1}t_{1}$, $s_{4}t_{2}$, $us_{1}$, $s_{4}v$ and $s_{2}s_{3}$. See Figure 4 for an illustration.
Figure 4: A special edge between $u$ and $v$.
Lemma 22.
Suppose an instance of the NTU partitioned matching game contains a special edge between $u$ and $v$. This special edge cannot belong to a matching in the weak core, but it can belong to a strongly blocking coalition in the following sense:
  1. i)
    If there is a matching $M$ in the weak core, then $s_{2}a_{1}$, $s_{1}t_{1}$ and $s_{4}t_{2}$ has to belong to $M$. This means that $us_{1}$ and $s_{4}v$ cannot belong to $M$.
  2. ii)
    Suppose the edges $us_{1}$ and $s_{4}v$ do not belong to a matching $M$. If by replacing the special edge between $u$ and $v$ with an edge $uv$, the players containing $u$ and $v$ would belong to a strongly blocking coalition w.r.t. $M$ restricted to the graph we get if we delete the special edge, then $M$ does not belong to the weak core.
@@
i) We have seen that if there are no additional edges, then $H$ cannot admit a matching in the weak core. Therefore if $s_{2}a_{1}$ does not belong to $M$, then the whole instance could not admit a matching in the weak core.
Suppose $s_{1}t_{1}$ or $s_{4}t_{2}$ does not belong to $M$, so $t_{1}$ or $t_{2}$ is unmatched in $M$. We know that $s_{2}a_{1}$ belongs to $M$, therefore $s_{3}$ is unmatched in $M$. But this means that the players $S$ and $T$ with the matching $M^{\prime}=\{s_{1}t_{1}$, $s_{2}s_{3}$, $s_{4}t_{2}\}$ form a blocking coalition, because $M^{\prime}$ covers every vertex of $S\cup T$.
ii) We have already seen that if the edges $s_{1}t_{1}$, $s_{2}a_{1}$ and $s_{4}t_{2}$ do not belong to $M$, then $M$ is not in the weak core. Suppose these edges belong to $M$, so $s_{3}$ is unmatched in $M$. Let $M^{\prime}$ be the blocking matching. Let $M^{\prime\prime}$ be the matching we get if we leave out the edge $uv$ from $M^{\prime}$ and add the edges $us_{1}$, $s_{4}v$ and $s_{2}s_{3}$. The players that belonged to the strongly blocking coalition would also be better off by $M^{\prime\prime}$, so together with $S$, they form a strongly blocking coalition: $S$ prefers $M^{\prime\prime}$ because all vertices of $S$ are covered by $M^{\prime\prime}$. ∎
Proof of Theorem 21.
We reduce from 3-SAT. Given an instance $I$ of 3-SAT, we construct an instance $I^{\prime}$ of the NTU partitioned matching game. For every variable of $I$, we construct a variable gadget, and for every clause we construct a clause gadget. After that, we describe the interconnecting edges between the clause gadgets and the variable gadgets.
Variable gadget. For a variable $x_{i}$, we define four sets of vertices, $X_{i}$, $\bar{X_{i}}$, $Y_{i}$ and $\bar{Y_{i}}$. $X_{i}$ contains a vertex for every occurrence of the variable $x_{i}$ in unnegated form, and the vertices of $X_{i}$ belong to one player. $\bar{X_{i}}$ contains a vertex for every occurrence of the variable $x_{i}$ in negated form, and the vertices of $\bar{X_{i}}$ belong to one player. For every vertex in $X_{i}$, there is a separate vertex in $Y_{i}$ so that there is an edge between them, and each vertex of $Y_{i}$ belongs to a separate player of size one. Similarly, for every vertex in $\bar{X_{i}}$, there is a separate vertex in $\bar{Y_{i}}$ so that there is an edge between them, and each vertex of $\bar{Y_{i}}$ belongs to a separate player of size one. Every vertex of $Y_{i}$ is connected to every vertex of $\bar{Y_{i}}$ by a special edge. See Figure 5.
Figure 5: The variable gadget. The dotted lines represent special edges.
@@
Interconnecting edges Now we describe the edges between the clause gadgets and the variable gadgets. To every literal that appears in a clause, there is a corresponding vertex in the clause gadget and a corresponding vertex in the variable gadget. We connect these two vertices with an edge.
@@
For any matching $M$ in the weak core of $I^{\prime}$, there is a truth assignment satisfying $I$.
@@
If $M$ is in the weak core, then $Y_{i}$ or $\bar{Y_{i}}$ has all its vertices covered by $M$. Indeed, if there were a vertex $u\in Y_{i}$ and a vertex $v\in\bar{Y_{i}}$ so that neither is covered by $M$, then the players $\{u\}$ and $\{v\}$ together with the special edge connecting them would form a strongly blocking coalition (see Lemma 22). Since the special edges cannot belong to $M$, every vertex of $X_{i}$ (or $\bar{X_{i}}$) is matched to a vertex in $Y_{i}$ (or $\bar{Y_{i}}$). In the first case we set the variable $x_{i}$ to be false, and in the second case we set it to be true. If both hold simultaneously, we arbitrarily set $x_{i}$ to be true or false, but it still holds that if $x_{i}$ is set to false, every vertex in $X_{i}$ is matched to a vertex in $Y_{i}$ in $M$.
@@
For any truth assignment satisfying $I$, it is possible to construct a matching $M$ in the weak core of $I^{\prime}$.
@@
If a variable $x_{i}$ is set to true (false), let $M$ contain all the edges between $\bar{X_{i}}$ ($X_{i}$) and $\bar{Y_{i}}$ ($Y_{i}$), and all interconnecting edges incident to $X_{i}$ ($\bar{X_{i}}$), see Figure 6.
Figure 6: Red edges are in $M$. The dotted lines represent special edges.
In a special edge $uv$, let $s_{1}t_{1}$, $s_{2}a_{1}$, $s_{4}t_{2}$ and $c_{2}c_{3}$, $b_{4}b_{5}$, $a_{2}a_{3}$, $c_{4}c_{5}$, $b_{2}b_{3}$, $a_{4}a_{5}$, $b_{6}c_{6}$, $b_{7}c_{7}$ belong to $M$, see Figure 7.
Figure 7: The red edges belong to $M$.
In every clause gadget, there is at least one vertex corresponding to a literal that is matched via an interconnecting edge in $M$ since every clause contains a literal set to true. Suppose that $a_{1}^{j}$ is such a vertex in the clause gadget corresponding to $c_{j}$. Let the edges $c_{2}^{j}c_{3}^{j}$, $b_{4}^{j}b_{5}^{j}$, $a_{2}^{j}a_{3}^{j}$, $c_{4}^{j}c_{5}^{j}$, $b_{2}^{j}b_{3}^{j}$, $a_{4}^{j}a_{5}^{j}$, $b_{6}^{j}c_{6}^{j}$ and $b_{7}^{j}c_{7}^{j}$ belong to $M$.
Now we show for every player that it cannot belong to a strongly blocking coalition. If the variable $x_{i}$ is set to true, then all the vertices in $X_{i}$, $\bar{X_{i}}$ and $\bar{Y_{i}}$ are matched in $M$, therefore the players $X_{i}$, $\bar{X_{i}}$ and the singleton players of $\bar{Y_{i}}$ cannot belong to a strongly blocking coalition. Since a vertex (which is also a player) in $Y_{i}$ is only connected to one of these players, it also cannot belong to a blocking coalition. If $x_{i}$ is set to false, a similar argument shows that these players cannot belong to a strongly blocking coalition.
Next, we show that a player of a special edge cannot belong to a strongly blocking coalition. $T$ cannot, since it is fully matched in $M$. $S$ cannot, since it has only one vertex unmatched in $M$, and $s_{1}$ is only connected to a vertex in $Y_{i}$ or $\bar{Y_{i}}$ and a vertex in $T$, therefore it would be unmatched in any blocking coalition. It remains to check that some of the players $A$, $B$ and $C$ do not form a blocking coalition. This is because the maximum matchings in $A$, $B$, $C$, $A\cup B\cup C$ and $B\cup C$ cover at most as many vertices as $M$, and a maximum matching in $A\cup B$ (or $A\cup C$) covers 12 vertices while $M$ covers 11 vertices of $A\cup B$ (and $A\cup C$).
This last argument about the players of the example shows that the players of a clause gadget also cannot belong to a strongly blocking coalition, which completes the proof. ∎
@@ 4.2 Strong core
It is coNP-complete to decide if a matching $M$ is in the strong core, even if each partition class has size at most 3. Furthermore, it is NP-hard to decide if the strong core is empty, even if each partition class has size at most 3. Both hardness results hold for bipartite graphs.
@@
We reduce from the NP-complete problem x3c as in the proof of Theorem 17; we use the same notation as in that proof. Let $I$ be an instance of x3c. We create an instance $I^{\prime}$ of the international matching game as follows. First we describe the vertex set $V$ of the graph.
@@
  • We also create connector vertices $b_{1,2}^{1},b_{1,2}^{2},b_{2,3}^{2},b_{2,3}^{3}\dots,b_{3n-2,3n-1}^{3n-2},b_{3n-2,3n-1}^{3n-1}$ and $y_{1,2}^{1},y_{1,2}^{2},y_{2,3}^{2},$ $y_{2,3}^{3},\dots,y_{3n-2,3n-1}^{3n-2},y_{3n-2,3n-1}^{3n-1}$.
@@
  • For each $i\in[3n-1]$ we have an edge $x_{i}a_{i}$ (note that there is no $x_{3n}$ vertex).
@@
  • For each set $S_{j}=\{j_{1},j_{2},j_{3}\}$ with $j_{1}<j_{2}<j_{3}$ we have edges $c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j}^{3}$.
The graph is bipartite with bipratition $U=\{c_{1}^{1},c_{1}^{2},c_{1}^{3},\dots,c_{m}^{1},c_{m}^{2},c_{m}^{2},x_{1},\dots,x_{3n-1}\}\cup\{b_{i,i+1}^{i},y_{i,i+1}^{i}\mid i\in[3n-1]\}$ and $W=V\setminus U$.
Finally, we describe the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2},d_{j}^{3}\}$. For each $i\in\{2,\dots,3n-2\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$ and a player $X_{i}=\{y_{i-1,i}^{i},x_{i},y_{i,i+1}^{i}\}$ Finally, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, $X_{1}=\{x_{1},y_{1,2}^{1}\}$ and $A_{3n-1}=\{b_{3n-2,3n-1}^{3n-1},a_{3n-1},a_{3n}\}$, $X_{3n-1}=\{y_{3n-2,3n-1}^{3n-1},x_{3n-1}\}$. All players have size at most 3.
We start by showing that if there is a matching $M$ in the strong core of $I^{\prime}$, then $M$ is the matching that contains all edges of $E$, except the ones of the form $c_{j}^{l}a_{j_{l}}$. First, observe that $M$ must contain all edges $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[3]$. Suppose that there is an edge $c_{j}^{l}d_{j}^{l}$ not in $M$. Then, player $D_{j}$ does not have all its vertices covered in $M$. However, $C_{j}$ and $D_{j}$ can create a matching among themselves that matches each of their vertices, so $\{C_{j},D_{,}\}$ would be a weakly blocking coalition to $M$, a contradiction. This implies that none of the edges of the form $c_{j}^{l}a_{j_{l}}$ are in $M$, as the $c_{j}^{l}$ vertices are matched elsewhere. Finally, observe that all the remaining edges can be added simultaneously, hence $M$ must contain all of them, as any strong core matching is maximal. Note that this matching $M$ matches every vertex of $V$, except $a_{3n}$.
Therefore, by showing that this matching $M$ is in the strong core if and only if there is no exact 3-cover, we show the desired hardness of both problems.
@@
If there is an exact 3-cover in $I$, then $M$ is not in the strong core.
@@
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $C_{l_{1}},\dots,C_{l_{n}},A_{1},A_{2},\dots,A_{3n-1}$ form a weakly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i=1,\dots,3n-2$. Hence, there is a matching $M^{\prime}$ among these players that matches every single one of their vertices. As $A_{3n-1}$ had a vertex exposed in $M$, it follows that $M$ is not in the strong core. ∎
@@
If $M$ is not in the strong core, then there is an exact 3-cover in $I$.
@@
Now, we make two observations. The first is that none of the players $X_{1},\dots,X_{3n-1}$ can participate in $\mathcal{P}$. Suppose for the contrary that $X_{i}\in\mathcal{P}$. Then, $M^{\prime}$ covers $y_{i-1,i}^{i},y_{i,i+1}^{i}$ (whenever such a vertex exists). Hence, $X_{i-1},X_{i+1}\in\mathcal{P}$ and by iterating this argument we get that $X_{i}\in\mathcal{P}$ for $i\in[3n-1]$. However, $a_{j_{1}}$ is not matched to $x_{j_{1}}$, hence $x_{j_{1}}$ is exposed in $M^{\prime}$, so player $X_{j_{1}}$ does not weakly improve, contradiction. The second observation is that players $A_{1},\dots,A_{3n-1}$ are all in $\mathcal{P}$. We have already seen that $A_{3n-1}\in\mathcal{P}$. Therefore, to weakly improve, $b_{3n-2,3n-1}^{3n-1}$ is matched, so $b_{3n-2,3n-1}^{3n-1}b_{3n-2,3n-1}^{3n-2}\in M^{\prime}$. This implies that $A_{3n-2}\in\mathcal{P}$ and by similar arguments we get that $A_{3n-3},\dots,A_{1}$ are in $\mathcal{P}$.

### v2

@@ 4.1 Weak core
We use Example 27 in our reductions to prove that deciding the emptiness of the weak core is hard.
The NO-gadget.
We introduce the key gadget in the proof, the NO-gadget. A NO-gadget between $u$ and $v$ will ensure that (i) neither $u$, nor $v$ can be connected to the NO-gadget in a matching in the weak core, but (ii) if a matching satisfies (i) for every NO-gadget and contains appropriate edges from each NO-gadget, then the matching is in the weak core if and only if it is in the weak core of the instance obtained by substituting each NO-gadget (between some vertices $u^{\prime},v^{\prime}$) by an edge ($u^{\prime}v^{\prime}$). Thus, we can use NO-gadgets to simulate forbidden edges.
Specifically, a NO-gadget between $u$ and $v$ is constructed as follows. Let $H_{uv}$ be a copy of the instance in Example 27 with players $A_{H_{uv}},B_{H_{uv}},C_{H_{uv}}$ and let $a_{1}^{H_{uv}}$ be a vertex of $A_{H_{uv}}$ in the clique of size 5 which contains only a single vertex of $A_{H}$. We add a player $S^{uv}$ with four vertices $s_{1}^{uv},s_{2}^{uv},s_{3}^{uv}$ and $s_{4}^{uv}$, and a player $T^{uv}$ with two vertices $t_{1}^{uv}$ and $t_{2}^{uv}$.
The edges are the following. We keep all edges of $H_{uv}$ and add $us_{1}^{uv},s_{1}^{uv}t_{1}^{uv},s_{2}^{uv}s_{3}^{uv},s_{2}^{uv}a_{1}^{H_{uv}},s_{4}^{uv}t_{2}^{uv},s_{4}^{uv}v$. See Figure 6 for an illustration.
Figure 6: The NO-gadget between $u$ and $v$. Upper indices have been omitted, since the figure illustrates only a single NO-gadget.
Let NTU-PMG-WC-Exist-F denote the problem, which asks if there is a matching in the weak core that excludes a set $F$ of forbidden edges.
Lemma 28.
NTU-PMG-WC-Exist-F with players of size at most $d$ can be reduced to NTU-PMG-WC-Exist with players of size at most $\max\{7,d\}$.
Proof.
Let $I=(G,\Pi,F)$ be an instance of NTU-PMG-WC-Exist-F with players of size at most $d$. We create an instance $I^{\prime}=(G^{\prime},\Pi^{\prime})$ of NTU-PMG-WC-Exist by replacing each edge $uv\in F$ with a NO-gadget between $u$ and $v$. As players in the NO-gadget have size at most 7, $I^{\prime}$ has players of size at most $\max\{d,7\}$.
We claim that $I$ admits a weak core matching disjoint from $F$ if and only if $I^{\prime}$ admits a matching in the weak core.
Suppose that $I^{\prime}$ has a matching $M^{\prime}$ in the weak core. Let $M=M^{\prime}\cap E(G)$. As we have seen in Example 27, for any NO-gadget, the copy $H_{uv}$ of Example 27 admits no weak core matching without additional edges, so $s_{2}^{uv}a_{1}^{H_{uv}}\in M^{\prime}$ for all NO-gadgets. Then, we also have that $s_{1}^{uv}t_{1}^{uv}\in M^{\prime}$ and $s_{4}^{uv}t_{2}^{uv}\in M^{\prime}$, else $\{T^{uv},S^{uv}\}$ is a strongly blocking coalition to $M^{\prime}$ with matching $\{s_{1}^{uv}t_{1}^{uv},s_{2}^{uv}s_{3}^{uv},s_{4}^{uv}t_{2}^{uv}\}$. By this observation, any edge of $M^{\prime}$ incident to a vertex $v\in V(G)$ is also present in $M$, so $|V_{i}\cap V(M)|=|V_{i}\cap V(M^{\prime})|$ for all $V_{i}\in\Pi$. Furthermore, it also follows that $M\cap F=\emptyset$.
Suppose for the contrary that $M$ is not in the weak core of $I$. Let $\mathcal{P}$ be a strongly blocking coalition to $M$ with some matching $M_{\mathcal{P}}$. We claim that $\mathcal{P}\cup\{S^{uv}\mid uv\in M_{\mathcal{P}}\cap F\}$ strongly blocks $M^{\prime}$ with the matching $M_{\mathcal{P}}\setminus F\cup\{us_{1}^{uv},s_{2}^{uv}s_{3}^{uv},s_{4}^{uv}v\mid uv\in M_{\mathcal{P}}\cap F\}$. Indeed, as $|V_{i}\cap V(M)|=|V_{i}\cap V(M^{\prime})|$ for all $V_{i}\in\Pi$, all players $i\in\mathcal{P}$ improve from $M^{\prime}$. Furthermore, the participating $S^{uv}$ players get all their vertices matched, while $s_{3}^{uv}$ was unmatched in $M^{\prime}$ (as $s_{1}^{uv}a_{1}^{H_{uv}}\in M^{\prime}$). This contradicts that $M^{\prime}$ is in the weak core.
Assume that $I$ admits a matching $M$ in the weak core such that $M\cap F=\emptyset$. Then, we can define a matching $M^{\prime}$ in $I^{\prime}$ as $M\cup\{s_{1}^{uv}t_{1}^{uv},s_{2}^{uv}a_{1}^{H_{uv}},s_{4}^{uv}t_{2}^{uv},b_{6}^{H_{uv}}c_{6}^{H_{uv}},b_{7}^{H_{uv}}c_{7}^{H_{uv}},a_{i}^{H_{uv}}a_{i+1}^{H_{uv}},b_{i}^{H_{uv}}b_{i+1}^{H_{uv}},c_{i}^{H_{uv}}c_{i+1}^{H_{uv}}\mid uv\in F,i=2,4\}$. See Figure 7 for an illustration.
We claim that $M^{\prime}$ is in the weak core of $I^{\prime}$. Suppose that $\mathcal{P}^{\prime}$ blocks $M^{\prime}$ with a matching $M^{\prime}_{\mathcal{P}^{\prime}}$. First, observe that no player $T^{uv}$ can be in $\mathcal{P}^{\prime}$, since both vertices are matched in $M^{\prime}$. Hence, if $S^{uv}\in\mathcal{P}^{\prime}$ for some $uv\in E(G)$, then all of its vertices must be matched, so for any $uv\in F$ either we have that $\{us_{1}^{H_{uv}},s_{4}^{H_{uv}}v\}\subseteq M^{\prime}_{\mathcal{P}^{\prime}}$ or $\{us_{1}^{H_{uv}},s_{4}^{H_{uv}}v\}\cap M^{\prime}_{\mathcal{P}^{\prime}}=\emptyset$.
Let $M_{\mathcal{P}}$ be the matching we get in $I$ by deleting any edge from a NO-gadget in $M^{\prime}_{\mathcal{P}^{\prime}}$, and adding the edge $uv$ if $\{us_{1}^{H_{uv}},s_{4}^{H_{uv}}v\}\subseteq M^{\prime}_{\mathcal{P}^{\prime}}$. Then, $M_{\mathcal{P}}$ is a matching and for any $V_{i}\in\Pi$, $|V_{i}\cap M_{\mathcal{P}}|=|V_{i}\cap M^{\prime}_{\mathcal{P}^{\prime}}|$. As $|V_{i}\cap M|=|V_{i}\cap M^{\prime}|$ also holds for $V_{i}\in\Pi$, this gives a contradiction, unless $M_{\mathcal{P}}=\emptyset$ and $\mathcal{P}^{\prime}$ contain only players in NO-gadgets. Then, no $S^{uv}$ player can be in $\mathcal{P}^{\prime}$ either, as $s_{1}^{uv},s_{4}^{uv}$ cannot be matched.
Hence, it only remains to show that the players of some copy $H_{uv}$ of Example 27 cannot form a strongly blocking coalition. It is easy to see that all three of them cannot form a strongly blocking coalition. Also, no player among them can match $6$ of their vertices alone to improve. Finally, for any two players, at least 2 of their vertices will remain unmatched in any matching, but at most $3$ of them are unmatched in $M^{\prime}$, so they cannot both improve. ∎
Figure 7: The edges in a NO-gadget in $M^{\prime}$ in the proof of Lemma 28. Bold edges belong to $M^{\prime}$.
@@
NTU-PMG-WC-Exist-7 is NP-hard.
@@
By Lemma 28, it is enough to show NP-hardness of NTU-PMG-WC-Exist-F, with players of size at most 7.
We reduce from 3-SAT, where we can assume that each variable appears at most four times. Given an instance $I$ of 3-SAT, we construct an instance $I^{\prime}$ of the NTU-PMG-WC-Exist-F. For every variable of $I$, we construct a variable gadget, and for every clause we construct a clause gadget. After that, we describe the interconnecting edges between the clause gadgets and the variable gadgets.
Variable gadget. For a variable $x_{i}$, we define four sets of vertices, $X_{i}$, $\bar{X_{i}}$, $Y_{i}$ and $\bar{Y_{i}}$. $X_{i}$ contains a vertex for every occurrence of the variable $x_{i}$ in unnegated form, and the vertices of $X_{i}$ belong to one player. $\bar{X_{i}}$ contains a vertex for every occurrence of the variable $x_{i}$ in negated form, and the vertices of $\bar{X_{i}}$ belong to one player. For every vertex in $X_{i}$, there is a separate vertex in $Y_{i}$ so that there is an edge between them, and each vertex of $Y_{i}$ belongs to a separate player of size one. Similarly, for every vertex in $\bar{X_{i}}$, there is a separate vertex in $\bar{Y_{i}}$ so that there is an edge between them, and each vertex of $\bar{Y_{i}}$ belongs to a separate player of size one. Every vertex of $Y_{i}$ is connected to every vertex of $\bar{Y_{i}}$ by a forbidden edge in $F$. See Figure 8.
Figure 8: The variable gadget for $x_{i}$. The dotted lines represent forbidden edges.
@@
Interconnecting edges. Now we describe the edges between the clause gadgets and the variable gadgets. To every literal that appears in a clause, there is a corresponding vertex in the clause gadget and a corresponding vertex in the variable gadget. We connect these two vertices with an edge.
As every variable appears at most four times, all players have size at most 7.
@@
For any matching $M$ in the weak core of $I^{\prime}$, there is a truth assignment that satisfies $I$.
@@
If $M$ is in the weak core, then $Y_{i}$ or $\bar{Y_{i}}$ has all its vertices covered by $M$. Indeed, if there were a vertex $u\in Y_{i}$ and a vertex $v\in\bar{Y_{i}}$ so that neither is covered by $M$, then the players $\{u\}$ and $\{v\}$ together with the forbidden edge connecting them would form a strongly blocking coalition (see Lemma 28). Since forbidden edges cannot belong to $M$, every vertex of $X_{i}$ (or $\bar{X_{i}}$) is matched to a vertex in $Y_{i}$ (or $\bar{Y_{i}}$). In the first case, we set the variable $x_{i}$ to be false, and in the second case we set it to be true. If both hold simultaneously, we arbitrarily set $x_{i}$ to be true or false, but it still holds that if $x_{i}$ is set to false, every vertex in $X_{i}$ is matched to a vertex in $Y_{i}$ in $M$.
@@
For any truth assignment satisfying $I$, it is possible to construct a matching $M$ in the weak core of $I^{\prime}$ without any forbidden edges.
@@
If a variable $x_{i}$ is set to true (false), let $M$ contain all the edges between $\bar{X_{i}}$ and $\bar{Y_{i}}$ (between $X_{i}$ and $Y_{i}$), and all interconnecting edges incident to $X_{i}$ ($\bar{X_{i}}$), see Figure 9.
Figure 9: Matching edges in the variable gadget when $x_{i}$ is set to true. Bold edges are in $M$. The dotted lines represent special edges.
In every clause gadget, there is at least one vertex corresponding to a literal that is matched via an interconnecting edge in $M$ since every clause contains a literal set to true. Suppose that $a_{1}^{j}$ is such a vertex in the clause gadget corresponding to $c_{j}$ (this does not forbid that $b_{1}^{j}$ and $c_{1}^{j}$ are such vertices too). Let the edges $c_{2}^{j}c_{3}^{j}$, $b_{4}^{j}b_{5}^{j}$, $a_{2}^{j}a_{3}^{j}$, $c_{4}^{j}c_{5}^{j}$, $b_{2}^{j}b_{3}^{j}$, $a_{4}^{j}a_{5}^{j}$, $b_{6}^{j}c_{6}^{j}$ and $b_{7}^{j}c_{7}^{j}$ belong to $M$. The other two cases are symmetric.
Then, $M$ contains no forbidden edges from $F$.
Now we show for every player that it cannot belong to a strongly blocking coalition. If the variable $x_{i}$ is set to true, then all the vertices in $X_{i}$, $\bar{X_{i}}$ and $\bar{Y_{i}}$ are matched in $M$; therefore, players $X_{i}$, $\bar{X_{i}}$ and the singleton players of $\bar{Y_{i}}$ cannot belong to a strongly blocking coalition. Since a vertex (which is also a player) in $Y_{i}$ is only connected to one of these players, it also cannot belong to a blocking coalition. If $x_{i}$ is set to false, a similar argument shows that these players cannot belong to a strongly blocking coalition.
Finally, similarly as in the proof of Lemma 28, the players in the clause gadgets also cannot form a strongly blocking coalition by themselves, as at most one of them has 2 vertices unmatched, the rest have at most one. ∎
@@ 4.2 Strong core
NTU-PMG-SC-Verif-3 is coNP-complete and NTU-PMG-SC-Exist-3 is coNP-hard. Both hardness results hold for bipartite graphs.
@@
We reduce from the NP-complete problem x3c as in the proof of Theorem 24; we use the same notation as in that proof. Let $I$ be an instance of x3c. We create an instance $I^{\prime}=(G,\Pi)$ of NTU-PMG-SC-Exist and an instance $I^{\prime\prime}=(G,\Pi,M)$ of NTU-PMG-SC-Verif as follows. Note that $G$ and $\Pi$ will be the same in $I^{\prime}$ and $I^{\prime\prime}$.
First, we describe the vertex set $V$ of the $G$.
@@
  • We create connector vertices $b_{1,2}^{1},b_{1,2}^{2},b_{2,3}^{2},b_{2,3}^{3}\dots,b_{3n-2,3n-1}^{3n-2},b_{3n-2,3n-1}^{3n-1}$ and $y_{1,2}^{1},y_{1,2}^{2},y_{2,3}^{2},y_{2,3}^{3},\dots$, $y_{3n-2,3n-1}^{3n-2},y_{3n-2,3n-1}^{3n-1}$.
@@
  • For each $i\in[3n-1]$ we have an edge $x_{i}a_{i}$ (note that there is no vertex $x_{3n}$).
@@
The graph is bipartite with bipartition $U=\{c_{1}^{1},c_{1}^{2},c_{1}^{3},\dots,c_{m}^{1},c_{m}^{2},c_{m}^{2},x_{1},\dots,x_{3n-1}\}\cup\{b_{i,i+1}^{i},y_{i,i+1}^{i}\mid i\in[3n-1]\}$ and $W=V\setminus U$.
Finally, we describe the partition $\Pi$ of the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2},d_{j}^{3}\}$. For each $i\in\{2,\dots,3n-2\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$ and a player $X_{i}=\{y_{i-1,i}^{i},x_{i},y_{i,i+1}^{i}\}$ Finally, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, $X_{1}=\{x_{1},y_{1,2}^{1}\}$ and $A_{3n-1}=\{b_{3n-2,3n-1}^{3n-1},a_{3n-1},a_{3n}\}$, $X_{3n-1}=\{y_{3n-2,3n-1}^{3n-1},x_{3n-1}\}$. All players have size at most 3.
We start by showing that if there is a matching $M^{\prime}$ in the strong core of $I^{\prime}$, then $M^{\prime}$ is the matching that contains all edges of $E$, except the ones of the form $c_{j}^{l}a_{j_{l}}$. First, observe that $M^{\prime}$ must contain all edges $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[3]$. Suppose that there is an edge $c_{j}^{l}d_{j}^{l}$ not in $M^{\prime}$. Then, player $D_{j}$ does not have all of its vertices covered in $M^{\prime}$. However, $C_{j}$ and $D_{j}$ can create a matching among themselves that matches each of their vertices, so $\{C_{j},D_{,}\}$ would be a weakly blocking coalition to $M^{\prime}$, a contradiction. This implies that none of the edges of the form $c_{j}^{l}a_{j_{l}}$ are in $M^{\prime}$, as the $c_{j}^{l}$ vertices are matched elsewhere. Finally, observe that all the remaining edges can be added simultaneously, hence $M^{\prime}$ must contain all of them, as any strong core matching is (inclusionwise) maximal. Note that this matching $M^{\prime}$ matches every vertex of $V$, except $a_{3n}$.
We let this matching be the matching $M$ in $I^{\prime\prime}$. Therefore, by showing that the matching $M$ is in the strong core of $I^{\prime\prime}$ if and only if there is no exact 3-cover in $I$, we also show that the strong core is nonempty in $I^{\prime}$ if and only if there is no exact 3-cover in $I$.
@@
If there is an exact 3-cover in $I$, then $M$ is not in the strong core in $I^{\prime\prime}$.
@@
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $C_{l_{1}},\dots,C_{l_{n}},A_{1},A_{2},\dots,A_{3n-1}$ form a weakly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i=1,\dots,3n-2$. Hence, there is a perfect matching on the vertex set of these players. As $A_{3n-1}$ had a vertex exposed in $M$, it follows that $M$ is not in the strong core. ∎
@@
If $M$ is not in the strong core of $I^{\prime\prime}$, then there is an exact 3-cover in $I$.
@@
Now, we make two observations. The first is that none of the players $X_{1},\dots,X_{3n-1}$ can participate in $\mathcal{P}$. Suppose for the contrary that $X_{i}\in\mathcal{P}$. Then, $M^{\prime}$ covers $y_{i-1,i}^{i},y_{i,i+1}^{i}$ (whenever such a vertex exists). Hence, $X_{i-1},X_{i+1}\in\mathcal{P}$ and by iterating this argument we get that $X_{i}\in\mathcal{P}$ for $i\in[3n-1]$. However, $a_{j_{1}}$ is not matched to $x_{j_{1}}$, hence $x_{j_{1}}$ is exposed in $M^{\prime}$, so player $X_{j_{1}}$ does not weakly improve, a contradiction. The second observation is that players $A_{1},\dots,A_{3n-1}$ are all in $\mathcal{P}$. We have already seen that $A_{3n-1}\in\mathcal{P}$. Therefore, to weakly improve, $b_{3n-2,3n-1}^{3n-1}$ is matched, so $b_{3n-2,3n-1}^{3n-1}b_{3n-2,3n-1}^{3n-2}\in M^{\prime}$. This implies that $A_{3n-2}\in\mathcal{P}$ and by similar arguments we get that $A_{3n-3},\dots,A_{1}$ are in $\mathcal{P}$.